0 0
Add a comment Improve this question Transcribed image text
Answer #1

-------------------------------------------------
Answer #2

problem 1)

since cement bag is at rest, net horizontal and vertical forces acting are zero

equating vertical force

Fg = 525 N = T3

T3 = T1*sin(65) +T2*sin(41)

525 = T1*sin(65) +T2*sin(41)............(1)

equating horizontal forces

T1*cos(65) = T2*cos(41)

T1 = 1.78 * T2................(2)

sustitutung T1 from (2) in....(1)

525 = 1.78*T2*sin(65) + T2*sin(45)

525 = 2.32*T2

T2 = 226.3 N

T1 = 1.78*226.3 = 402.814 N

T3 = 525 N

Problem 2)Using the kinematic equation

a)d = Vi*t + (1/2)*a*t^2

Vi = initial velocity = 0

d=1.9 m

1.9 = (1/2)*a*1.8^2

acceleration =a = 1.173 m/s^2

B)force on the block along the inclination

F = ma = mgcos(60) - k*mgsin(60)

a = 9.8*(0.5) - k*9.8*0.866

1.173 - 4.9 = -k*8.48

k = 4.395 (coeeficient of kinetic fricition)

C) Force of friction = k * mgsin(60) = 96.915 N

direction is upwards along the inclination

D) Vf = Vi +at

Vi = 0 m/s

Vf = 1.173*1.8 = 2.1114 m/s